OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x. x f(x) = 9x + 7 g(x) = 5x ^x -2 -11 -1 -2 0 1 1 5 2 Complete the chart and determine the solution of the equation f(x) = g(x)? x = -1 x = 0 x = 2 x = 25

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Screen shot it please

OpenStudy (anonymous):

how do you screen shot

OpenStudy (the_fizicx99):

go to start tab, search snipping tool then save it to pictures and click the "Attach file" on the bottom left

OpenStudy (the_fizicx99):

x = 2 :)

OpenStudy (anonymous):

thank you so much do you HAVE TIME FOR ONE MORE

OpenStudy (agent0smith):

g(x) = 5x ^x is clearly not the same as g(x) = 5^x btw

OpenStudy (the_fizicx99):

Agent will put my work to shame so I'll try to make it look good xD f(2) = 9(2) + 7 f(2) = 18 + 7 f(2) = 25 It's 5^x, 5^2 = 25

OpenStudy (anonymous):

oh sorry for the confusion

OpenStudy (the_fizicx99):

This is FLVS! O.O

OpenStudy (anonymous):

this is connexus

OpenStudy (the_fizicx99):

=.= I recognize them ._. wtf

OpenStudy (anonymous):

I think connexus and flvs are partners

OpenStudy (the_fizicx99):

._.